### v3

@@ Acknowledgements
This research is supported by National Natural Science Foundation of China (No. 62071127, and 62101137), National Key Research and Development Program of China (No. 2022ZD0160100), Shanghai Natural Science Foundation (No. 23ZR1402900) and Shanghai Municipal Foundation (No.2021SHZDZX0103). We acknowledge the ECMWF and NCEI for their great efforts in storing and providing the invaluable data, which is very important for this work and the research community. We would also like to express our appreciation to Mr. Huihang Sun and Prof. Jiayuan Fan for their suggestion and valuable discussions during the conduction of this research, and the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure.
Declarations
[…]
Competing interests The authors declare no competing interests.
[…]
Data availability The ERA5 dataset can be downloaded from the official website of Climate Data Store (CDS) at https://cds.climate.copernicus.eu/. The GDAS observational bufr files can be obtained from the Archive Information Request System (AIRS) of NCEI at https://www.ncei.noaa.gov/has/HAS.DsSelect.
[…]
Code availability The neural network model is developed using standard libraries in open-source platforms including PyTorch. Codes and model checkpoints used in this study will be available at https://github.com/kunc3301/FengWu-Adas if ready.
[…]
Author contribution L.B. and T.C. are co-corresponding authors. L.B., T.C., J.-J.L. and W.O. supervised this research. Kun Chen, L.B., F.L. and P.Y. conceived and initiated this project. Kun Chen and L.B. designed the details of model and developed the methodologies. Kun Chen and Kang Chen developed the Python code. Kun Chen, Y.X. and T.H. prepared and processed the data. Kun Chen performed experiments. Kun Chen, L.B., F.L., P.Y. and H.C. wrote the manuscript. All authors contributed to analysis of data, discussions about the results, and improvement of the presentation.
    Appendix A Extended Data
    Figure 7: RMSE and Bias skill of FengWu-Adas with random masks at each time step in simulation experiments. Rows 1 and 3 show the RMSE of the analyses for main variables over the year, and rows 2 and 4 show the Bias for them. There are no obvious performance differences compared to Fig. 3, which prove the assimilation ability and robustness of Adas. See the caption of Fig. 3 for all other details.
    Figure 8: Average RMSE skill of FengWu-Adas for all variables in simulation experiments. The different colors represent the RMSE of analyses with different observation ratios, as marked in the legend. The first subplot includes four surface variables, and other subplots correspond to five upper-air variables, as indicated in the subplot titles. The x-axis represents the single variable (and pressure level). The y-axis represents the RMSE.
    Figure 9: RMSE skill of FengWu-Adas with random initialization in simulation experiments. The different colors represent the results with different observation ratios, as marked in the bottom. Each subplot represents a single variable (and pressure level), as indicated in the subplot titles, and the subgraph provides a clearer display of the first ten days. The x-axis represents the operation time at 6-hour steps over the year. The y-axis represents the RMSE, which uses a coordinate axis with non equidistant logarithmic form to ensure visual perception. The results further demonstrate the ability of Adas to propagate sparse observation information and the stability of the system.
    Figure 10: RMSE and ACC skill of FengWu-Adas for weather forecasting over ten days in simulation experiments. The different colors represent the forecasts whose initial states are generated after assimilating observations with different ratios, as marked in the bottom. Based on the analysis produced by Adas, FengWu can still generate accurate and consistent global medium-range weather forecasts. See the caption of Fig. 6 for all other details.
    Figure 11: Average RMSE skill of FengWu-Adas for all variables in GDAS experiments. The different colors represent the RMSE of analyses after assimilating all data and non-satellite data, as marked in the legend. See the caption of Extended Data Fig. 8 for all other details.

### v4

@@ Acknowledgements
This research is supported by National Natural Science Foundation of China (No.62071127 and 62101137), National Key Research and Development Program of China (No.2022ZD0160101), Shanghai Natural Science Foundation (No.23ZR1402900), Shanghai Science and Technology Commission Explorer Program Project (24TS1401300) and Shanghai Municipal Science and Technology Major Project (No.2021SHZDZX0103). We acknowledge the ECMWF and NCEI for their great efforts in storing and providing the invaluable data, which are very important for this work and the research community. We would also like to express our appreciation to Dr. Ben Fei, Mr. Huihang Sun and Prof. Jiayuan Fan for their suggestions and valuable discussions during the conduct of this research, and the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure.
Data availability
The ERA5 dataset can be downloaded from the official website of Climate Data Store (CDS) at https://cds.climate.copernicus.eu. The GDAS observational bufr files can be obtained from the Archive Information Request System (AIRS) of NCEI at https://www.ncei.noaa.gov/has/HAS.DsSelect. The IGRA dataset and its documentation are publicly available at https://www.ncei.noaa.gov/data/integrated-global-radiosonde-archive.
Code availability
The neural network model is developed using standard libraries in open-source platforms including PyTorch. Codes and model checkpoints used in this study are available at https://github.com/OpenEarthLab/FengWu-Adas.